# MMSI-Bench: A Benchmark for Multi-Image Spatial Intelligence

### v1

@@ Abstract
Spatial intelligence is essential for multimodal large language models (MLLMs) operating in the complex physical world. Existing benchmarks, however, probe only single-image relations and thus fail to assess the multi-image spatial reasoning that real-world deployments demand. We introduce MMSI-Bench, a VQA benchmark dedicated to multi-image spatial intelligence. Six 3D-vision researchers spent more than 300 hours meticulously crafting 1,000 challenging, unambiguous multiple-choice questions from over 120,000 images, each paired with carefully designed distractors and a step-by-step reasoning process. We conduct extensive experiments and thoroughly evaluate 34 open-source and proprietary MLLMs, observing a wide gap: the strongest open-source model attains roughly 30% accuracy and OpenAI’s o3 reasoning model reaches 40%, while humans score 97%. These results underscore the challenging nature of MMSI-Bench and the substantial headroom for future research. Leveraging the annotated reasoning processes, we also provide an automated error analysis pipeline that diagnoses four dominant failure modes, including (1) grounding errors, (2) overlap-matching and scene-reconstruction errors, (3) situation-transformation reasoning errors, and (4) spatial-logic errors, offering valuable insights for advancing multi-image spatial intelligence.
Figure 1: The MMSI-Bench benchmark. Left and middle: Human experts carefully select sets of images that enable multi-image spatial reasoning to answer unambiguous and challenging spatial questions; human-annotated reasoning processes further facilitate automated error analysis. Right: Comparison of models and human performance on MMSI-Bench, highlighting the benchmark’s difficulty and the substantial gap between current MLLMs and human-level spatial intelligence.
@@ 1 Introduction
Multimodal large language models (MLLMs) have rapidly advanced their ability to ground language and vision in the physical world [35, 12, 47], making them a promising stepping-stone toward embodied artificial general intelligence (AGI). A core capability of MLLMs deployed in the physical world is spatial intelligence—the skill to understand where things are, how they move, etc. To steer progress toward this goal, the community needs rigorous, diverse, and challenging evaluation benchmarks that can serve as north stars for progress.
@@
However, most existing benchmarks probe only single-image spatial understanding and restrict themselves to straightforward spatial relations [34, 56, 27, 17, 15]. Real-world spatial understanding is more complex: models must reason across multiple images to track object and ego motion, and relate entities that never co-occur in a single frame [62, 58, 38]. Without benchmarks that demand such multi-image reasoning, we cannot reliably measure or improve an MLLM’s spatial competence.
@@
In this work, we introduce MMSI-Bench, a multiple-choice VQA benchmark dedicated to Multi-iMage-based Spatial Intelligence (Figure 1). MMSI-Bench is built on a comprehensive taxonomy of ten fundamental spatial reasoning tasks that span the positions, attributes, and motions of three key elements—camera, object, and region in real-world scenes. Beyond these atomic tasks, we construct a multi-step reasoning split that chains them into long-horizon questions.
@@
Crafting a benchmark that is simultaneously diverse, accurate, and challenging is far from trivial. Prior benchmarks [15, 17, 65, 22] populate templated questions from annotated or estimated spatial metadata. These automatic pipelines suffer from limited diversity and incomplete coverage. We therefore adopt a fully human-centric design. For each question, one of six 3D-vision researchers scours large image collections, selects a set of images that exhibit non-trivial spatial relations, and creates a novel, difficult, and unambiguous multiple-choice question whose answer can be derived only by reasoning across all selected images. Every item comes with carefully designed distractors and a step-by-step reasoning process that explicitly leads to the correct answer; a second annotator then audits each example to ensure clarity and correctness.
@@
Over the course of 300+ researcher-hours, we inspected more than 120,000 candidate images and distilled them into 1,000 high-quality Q&A pairs covering a broad spectrum of real-world scenes—from indoor scans [18, 14] and outdoor driving footage [54, 13] to robot manipulation [9] and everyday activities [23]. We evaluate 34 widely used MLLMs on MMSI-Bench. The strongest open-source model attains about 30% accuracy, while the best proprietary system, the OpenAI o3 model [46] with a reasoning mode, reaches only 40%, compared with human performance at 97% accuracy. To the best of our knowledge, MMSI-Bench shows the largest performance margin between state-of-the-art models and humans among existing spatial intelligence benchmarks.
@@ 2 Related Work
Multi-image VQA benchmarks. Driven by the rapid progress of MLLMs, VQA benchmarks have shifted from single-image reasoning [25] to multi-image settings focusing on various skills. Visual Haystacks  [60] probes long-context retrieval, asking models to locate relevant content within large image collections. MIBench  [36] stresses cross-image reasoning that combines instruction following with knowledge retrieval. ReMI [28] broadens the scope to math, physics, and code across two to six images, blending GPT-generated items with human audits. BLINK [22] zeroes in on inter-image relationships, while MuirBench [57] adds robustness and diversity to multi-image reasoning. Existing benchmarks show that contemporary MLLMs still trail behind humans as either the number of images or the reasoning complexity grows [31]. In this work, instead of building a general benchmark, we focus on comprehensively evaluating MLLMs’ multi-image spatial intelligence.
Benchmarks for spatial intelligence of MLLMs. Recent progress in MLLMs has enabled them to potentially serve as the “brains” of embodied agents, e.g., the decision-making modules of autonomous vehicles [53] or robots [29, 12, 32], where reliable spatial intelligence is indispensable. To measure this capability, a variety of benchmarks have been proposed, but the majority assess only single-image spatial understanding [33, 66, 52, 34, 27, 20, 56, 49, 39]. Beyond a single view, associating information across multiple images is even more critical in practice, yet existing multi-image benchmarks remain fragmentary. BLINK [22] and MuriBench [57] contain just a few spatial sub-splits within otherwise general VQA suites, and thus lack comprehensive coverage. LEGO-Puzzles [55] probes multi-step reasoning in a synthetic LEGO-assembly setting, overlooking real-world scenes. Efforts such as UniQA-3D [68], MMIU [40], SAT [50], MultiSPA [63], and VSI-Bench [65] generate template-based or rule-based questions from existing annotations/metadata or simulators, which constrains question diversity. By contrast, our MMSI-Bench is fully human-curated, free from template limitations, contains challenging and diverse questions, and offers rigorous evaluation of MLLMs’ spatial intelligence in multi-image scenarios.
@@ 3 MMSI-Bench
In this section, we present MMSI-Bench, a comprehensive and challenging benchmark designed to evaluate the multi-image spatial reasoning capability of MLLMs. Specifically, we introduce the concept of multi-image spatial reasoning and outlines its associated sub-tasks in Section 3.1. Subsequently, in Section 3.2, we provide a detailed description of the benchmark construction process.
Figure 2: Representative MMSI-Bench samples from each category. Please zoom in to inspect image details. Questions and rationales are simplified for brevity; the complete versions appear in the supplementary material. Correct answers are highlighted in green.
@@ 3.1 Overview of MMSI-Bench
Humans possess the remarkable ability to infer spatial information about their environment by integrating visual inputs from different viewpoints or moments in time. We refer to this process as multi-image spatial reasoning, in which information from multiple images is combined to reconstruct the underlying scene and enable reasoning to answer complex spatial questions. To evaluate the multi-image spatial reasoning capability of multimodal large language models, we introduce MMSI-Bench.
To systematically evaluate multi-image spatial reasoning, we categorize tasks based on three fundamental spatial elements: camera (the agent that captures the scenes), object (entities other than the agent within the scenes), and region (semantic areas within the scenes, such as a kitchen). Fundamental spatial understanding requires not only recognizing the positional relationships among these elements but also their attributes and potential motion. Guided by this taxonomy, we define ten spatial reasoning task types and one multi-step spatial reasoning category, which flexibly combines these basic types. Examples for all task categories are provided in Figure 2. Notably, our questions are designed to be human-answerable; thus, we do not include questions about camera attributes (parameters). Additionally, since regions are generally static, we do not include motion-related tasks for regions. The specific categories are as follows:
  • Positional Relationship (camera-camera): Inferring the positional relationship between two camera viewpoints.
  • Positional Relationship (camera-object): Identifying the positional relationship between the camera and an object.
  • Positional Relationship (camera-region): Assessing the positional relationship between the camera and a semantic region (e.g., kitchen).
  • Positional Relationship (object-object): Inferring the positional relationship between objects within the scenes.
  • Positional Relationship (object-region): Determining the positional relationship between an object and a semantic region.
  • Positional Relationship (region-region): Identifying the positional relationship between different regions within a given scene.
  • Attribute – Measurement: Inferring geometric properties (e.g., length, size) of elements.
  • Attribute – Appearance: Inferring visual characteristics of elements, such as their shape.
  • Motion (camera): Reasoning about the camera’s movement, such as its movement direction.
  • Motion (object): Assessing the movement of objects within the scene.
  • Multi-Step Reasoning: Solving complex tasks that require multiple sequential steps of the spatial reasoning types described above.
In conclusion, MMSI-Bench consists of 1,000 multiple-choice question–answer pairs based on multiple images, each also accompanied by annotated reasoning processes explaining the answer. Every question is specifically designed such that it cannot be answered using a single image alone. Except for the Multi-Step Reasoning category, all other categories contain only two images per question. Detailed statistics on the proportion of questions in each category and other relevant information about MMSI-Bench are presented in Figure 4 and Table 1, respectively.
@@
Table 1: Dataset statistics.
@@ 3.2 Benchmark Construction Process
Data collection. To comprehensively evaluate the capability of MLLMs across a variety of real-world scenarios, it is essential to curate data from diverse and representative sources. To ensure the complexity and authenticity of the evaluation, we exclusively utilize images captured from real-world scenes. Accordingly, we leverage images from a diverse collection of real-world datasets: indoor 3D scene datasets (ScanNet [18] and Matterport3D [14]), a local scene reconstruction dataset (DTU [26]), autonomous driving datasets (nuScenes [13], Waymo [54]), a robotics dataset (AgiBot-World [9]), a video object segmentation dataset (DAVIS 2017 [48]), and an egocentric video dataset (Ego4D [23]). These diverse data sources enable the construction of various question-answer pairs, ensuring that our benchmark captures the complexity and diversity inherent in real-world environments.
@@
Question-answer design and reasoning process Annotation. Based on the collected datasets, six 3D Vision researchers annotate QA pairs spanning the above eleven distinct task categories (with each category allocated to all six annotators). Annotators are explicitly instructed to create a diverse set of questions that are as novel and as challenging as possible for MLLMs, aiming to push the boundaries of current model capabilities. At the same time, annotators are instructed to ensure that all questions are accurate, unambiguous, and answerable by humans. For each question, annotators begin by visually inspecting the image database to choose a set of relevant images. They then construct a question that cannot be answered based on any single image in isolation; rather, the question requires integrating information from multiple images in order to infer the actual scenario. To ensure diversity in question formats, no templates are used, and all questions are created in free form. Each question is designed as a four-option multiple-choice problem, with only one correct answer. The incorrect options are also carefully crafted to serve as plausible distractors. In addition, annotators provide detailed reasoning processes alongside each answer, which are used to facilitate filtering incorrect samples during quality control and to enable automated evaluation and analysis of MLLMs’ reasoning processes, as described in Section 5.2.
@@
Quality Control. To ensure the quality of the data, three additional reviewers—independent of the original annotators—systematically examine all data and remove any instances containing ambiguous questions (due to linguistic ambiguities or insufficient visual information), incorrect answers, or questions that can be answered using a single image or common sense alone. This rigorous quality control ensures our benchmark provides a reliable, high-quality evaluation.
Models Positional Relationship Attribute Motion MSR Avg. Cam.–Cam. Obj.–Obj. Reg.–Reg. Cam.–Obj. Obj.–Reg. Cam.–Reg. Meas. Appr. Cam. Obj. Proprietary o3 45.2 39.4 37.0 44.2 47.1 62.6 54.7 28.8 31.1 32.9 34.9 41.0 GPT-4.5 34.4 29.8 39.5 51.2 47.1 55.4 39.1 33.3 41.9 40.8 36.4 40.3 GPT-4.1 36.6 26.6 27.2 29.1 36.5 27.7 37.5 24.2 36.5 32.9 28.8 30.9 GPT-4o 34.4 24.5 23.5 19.8 37.6 27.7 32.8 31.8 35.1 36.8 30.8 30.3 Gemini-2.5-Pro 39.7 31.9 39.5 45.3 35.2 43.3 51.5 21.2 36.4 30.2 34.3 36.9 –Thinking 37.6 30.9 42.0 39.5 40.0 42.2 50.0 27.3 35.1 32.9 34.3 37.0 Claude-3.7-Sonnet 32.3 26.6 22.2 34.9 37.6 42.2 25.0 22.7 21.6 32.9 22.7 28.7 –Thinking 35.5 29.8 29.6 31.4 32.9 34.9 37.5 25.8 23.0 31.6 25.8 30.2 Doubao-1.5-pro 25.8 33.0 33.3 33.7 49.4 39.8 37.5 31.8 27.0 26.3 29.8 33.0 Seed1.5-VL 32.2 30.8 25.9 23.2 38.8 32.5 39.0 21.2 36.4 25.0 26.2 29.7 Open-source InternVL3-78B 34.4 23.4 32.1 12.8 37.6 26.5 37.5 19.7 28.4 31.6 29.3 28.5 InternVL2.5-78B 23.7 22.3 39.5 29.1 31.8 42.2 35.9 19.7 17.6 26.3 27.3 28.5 Qwen2.5-VL-72B 25.8 34.0 34.6 23.3 34.1 36.1 45.3 27.3 27.0 30.3 27.3 30.7 LLaVA-OneVision-72B 43.0 31.9 33.3 30.2 37.6 38.6 28.1 19.7 13.5 32.9 15.7 28.4 InternVL3-38B 21.5 20.2 33.3 23.3 35.3 25.3 39.1 21.2 16.2 31.6 25.8 26.3 InternVL2.5-38B 18.3 22.3 35.8 22.1 38.8 34.9 37.5 25.8 14.9 38.2 25.3 27.9 Qwen2.5-VL-32B 24.7 26.6 29.6 22.1 32.9 31.3 31.2 24.2 18.9 35.5 27.8 27.7 InternVL2.5-26B 24.7 19.1 29.6 33.7 31.8 37.3 35.9 30.3 10.8 31.6 26.8 28.0 NVILA-15B 30.1 39.4 28.4 36.0 38.8 20.5 29.7 31.8 18.9 35.5 27.8 30.5 InternVL3-14B 19.4 24.5 24.7 23.3 37.6 24.1 31.2 22.7 24.3 31.6 29.3 26.8 Llama-3.2-11B-Vision 25.8 30.8 32.0 25.6 21.2 25.9 20.3 19.7 25.6 28.9 19.2 25.4 InternVL3-9B 18.3 25.5 32.1 29.1 31.8 22.9 29.7 24.2 16.2 38.2 26.8 26.7 InternVL3-8B 25.8 31.9 37.0 25.6 35.3 28.9 23.4 24.2 16.2 32.9 14.6 25.7 InternVL2.5-8B 32.3 27.7 29.6 32.6 24.7 32.5 26.6 27.3 16.2 31.6 30.3 28.7 NVILA-8B 17.2 29.8 24.7 30.2 22.4 34.9 34.4 25.8 25.7 34.2 29.8 28.1 Qwen2.5-VL-7B 24.7 24.5 24.7 25.6 29.4 26.5 25.0 18.2 20.3 39.5 25.8 25.9 LLaVA-OneVision-7B 20.4 33.0 29.6 29.1 25.9 30.1 29.7 25.8 18.9 34.2 11.6 24.5 InternVL2.5-4B 31.2 23.4 21.0 31.4 34.1 25.3 23.4 24.2 13.5 31.6 26.8 26.3 Qwen2.5-VL-3B 26.9 27.7 30.9 29.1 28.2 34.9 31.2 16.7 17.6 27.6 23.2 26.5 InternVL3-2B 26.9 25.5 29.6 31.4 28.2 27.7 26.6 22.7 12.2 23.7 23.7 25.3 InternVL2.5-2B 28.0 27.7 24.7 37.2 29.4 36.1 43.8 15.2 21.6 31.6 26.8 29.0 InternVL3-1B 24.7 35.1 22.2 30.2 29.4 30.1 32.8 28.8 17.6 19.7 26.3 27.0 InternVL2.5-1B 23.7 26.6 24.7 25.6 31.8 25.3 31.2 30.3 17.6 25.0 26.3 26.1 DeepSeek-VL2 23.7 31.9 22.2 36.0 30.6 22.9 28.1 15.2 28.4 26.3 28.3 27.1 DeepSeek-VL2-Small 24.7 28.7 18.5 33.7 38.8 27.7 28.1 33.3 24.3 25.0 29.8 28.6 DeepSeek-VL2-Tiny 29.0 27.7 21.0 23.3 17.6 31.3 14.1 24.2 14.9 25.0 27.3 24.0 Baseline Blind GPT-4o 20.2 17.0 29.6 13.9 29.4 19.2 21.8 12.1 20.2 29.0 20.2 22.7 Random Guessing 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 Human Level 95.7 98.9 97.5 94.2 98.8 96.4 95.3 98.5 98.6 98.7 97.0 97.2
Table 2: Evaluation results for 34 MLLMs. For each category, the best-performing proprietary model and the best-performing open-source model are both indicated in bold.

### v2

@@ Abstract
Spatial intelligence is essential for multimodal large language models (MLLMs) operating in the complex physical world. Existing benchmarks, however, probe only single-image relations and thus fail to assess the multi-image spatial reasoning that real-world deployments demand. We introduce MMSI-Bench, a VQA benchmark dedicated to multi-image spatial intelligence. Six 3D-vision researchers spent more than 300 hours meticulously crafting 1,000 challenging, unambiguous multiple-choice questions from over 120,000 images, each paired with carefully designed distractors and a stepwise reasoning process. We conduct extensive experiments and evaluate 37 open-source and proprietary MLLMs, observing a wide gap: the strongest open-source model attains roughly 30% accuracy and OpenAI’s GPT-5 reasoning model reaches 40%, while humans score 97%. These results underscore the challenging nature of MMSI-Bench and the substantial headroom for future research. Leveraging the annotated reasoning processes, we also provide an automated error analysis pipeline that diagnoses four dominant failure modes, including (1) grounding errors, (2) overlap-matching and scene-reconstruction errors, (3) situation-transformation reasoning errors, and (4) spatial-logic errors, offering insights for advancing spatial intelligence.
Figure 1: The MMSI-Bench benchmark. Left and middle: Human experts carefully select image sets to create unambiguous and challenging multi-image spatial reasoning questions. Human-annotated reasoning processes are also provided for correctness verification and to further facilitate automated error analysis. Right: Comparison of models and human performance on MMSI-Bench, highlighting the benchmark’s difficulty and the substantial gap between current MLLMs and human-level spatial intelligence.
@@ 1 Introduction
Multimodal large language models (MLLMs) have rapidly advanced their ability to ground language and vision in the physical world (Liu et al., 2024a, Black et al., 2024, Peng et al., 2023), making them a promising stepping-stone toward embodied artificial general intelligence (AGI). A core capability of MLLMs deployed in the physical world is spatial intelligence, the skill to understand where things are, how they move, etc. To steer progress toward this goal, the community needs rigorous, diverse, and challenging evaluation benchmarks that can serve as north stars for progress.
@@
However, most existing benchmarks probe only single-image spatial understanding and restrict themselves to straightforward spatial relations (Liu et al., 2023, Tong et al., 2024, Johnson et al., 2017, Cheng et al., 2024, Chen et al., 2024a). Real-world spatial understanding is more complex: models must reason across multiple images to track object and ego motion, and relate entities that never co-occur in a single frame (Xu et al., 2024, Wang et al., 2024b, Lyu et al., 2024). Without benchmarks that demand such multi-image reasoning, we cannot reliably measure or improve an MLLM’s spatial competence.
@@
In this work, we introduce MMSI-Bench, a multiple-choice VQA benchmark dedicated to Multi-iMage-based Spatial Intelligence (Figure 1). MMSI-Bench is built on a comprehensive taxonomy of ten fundamental spatial reasoning tasks that span the positions, attributes, and motions of three key spatial elements—camera (observer), object, and region in real-world scenes. Beyond these atomic tasks, we construct a multi-step reasoning split that chains them into long-horizon questions.
@@
Crafting a benchmark that is simultaneously diverse, accurate, and challenging is far from trivial. Prior benchmarks (Chen et al., 2024a, Cheng et al., 2024, Yang et al., 2024, Fu et al., 2024) populate templated questions from annotated or estimated spatial metadata. These automatic pipelines suffer from limited diversity and difficulty. We therefore adopt a fully human-centric design. For each question, one of six 3D-vision researchers scours large image collections, selects a set of images that exhibit non-trivial spatial relations, and creates a novel, difficult, and unambiguous multiple-choice question whose answer can be derived only by reasoning across all selected images. Every item comes with carefully designed distractors and a step-by-step reasoning process that explicitly leads to the correct answer, a second annotator then audits each example to ensure clarity and correctness.
@@
Over the course of 300+ researcher-hours, we inspected more than 120,000 candidate images and distilled them into 1,000 high-quality Q&A pairs covering a broad spectrum of real-world scenes, from indoor scans (Dai et al., 2017, Chang et al., 2017) and outdoor driving footage (Sun et al., 2020, Caesar et al., 2020) to robot manipulation (AgiBot-World-Contributors et al., 2025) and everyday activities (Grauman et al., 2022). We evaluate 37 widely used MLLMs on MMSI-Bench. The strongest open-source model attains about 30% accuracy, while the best proprietary system, the OpenAI GPT-5 reasoning model (OpenAI, 2025c), reaches only 40%, compared with human performance at 97% accuracy. To the best of our knowledge, MMSI-Bench shows the largest performance margin between SOTA models and humans among existing spatial intelligence benchmarks.
@@ 2 Related Work
Multi-image VQA benchmarks. Driven by the rapid progress of MLLMs, VQA benchmarks have shifted from single-image reasoning (Hudson & Manning, 2019) to multi-image settings focusing on various skills. Visual Haystacks  (Wu et al., 2025b) probes long-context retrieval, asking models to locate relevant content within large image collections. MIBench  (Liu et al., 2024b) stresses cross-image reasoning that combines instruction following with knowledge retrieval. ReMI (Kazemi et al., 2024) broadens the scope to math, physics, and code across two to six images, blending GPT-generated items with human audits. BLINK (Fu et al., 2024) zeroes in on inter-image relationships, while MuirBench (Wang et al., 2024a) adds robustness and diversity to multi-image reasoning. Existing benchmarks show that contemporary MLLMs still trail behind humans as either the number of images or the reasoning complexity grows (Li et al., 2023). In this work, instead of building a general benchmark, we focus on comprehensively evaluating MLLMs’ multi-image spatial intelligence.
Benchmarks for spatial intelligence of MLLMs. Recent progress in MLLMs has enabled them to potentially serve as the “brains” of embodied agents, e.g., the decision-making modules of autonomous vehicles (Sima et al., 2024) or robots (Kim et al., 2024, Team et al., 2025a, Black et al., 2024), where reliable spatial intelligence is indispensable. To measure this capability, a variety of benchmarks have been proposed, but the majority are designed to assess only single-image spatial understanding (Liao et al., 2024, Liu et al., 2023, Shiri et al., 2024, Zhou et al., 2025, Ma et al., 2024, Song et al., 2024, Du et al., 2024, Tong et al., 2024, Zhang et al., 2024). Beyond a single view, associating information across multiple images is even more critical in practice, yet existing multi-image benchmarks remain fragmentary. BLINK (Fu et al., 2024) and MuriBench (Wang et al., 2024a) contain just a few spatial sub-splits within otherwise general VQA suites, and thus lack comprehensive coverage. LEGO-Puzzles (Tang et al., 2025) probes multi-step reasoning in a synthetic LEGO-assembly setting, overlooking real-world scenes. Efforts such as UniQA-3D (Zuo et al., 2024), MMIU (Meng et al., 2024), SAT (Ray et al., 2024), MultiSPA (Xu et al., 2025), and VSI-Bench (Yang et al., 2024) generate template-based or rule-based questions from existing annotations/metadata or simulators, which constrains question diversity and difficulty. By contrast, MMSI-Bench is fully human-curated, unconstrained by templates, and features challenging, diverse questions with accompanying reasoning annotations. ERQA (Team et al., 2025b) is likewise human-curated but comprises only 400 samples, of which just 113 are multi-image. Detailed quantitative comparisons (e.g., diversity) are provided in Appendix D.
@@ 3 MMSI-Bench
In this section, we present MMSI-Bench, a comprehensive and challenging benchmark designed to evaluate the multi-image spatial reasoning capability of MLLMs. Specifically, we introduce the concept of multi-image spatial reasoning and outline its associated sub-tasks in Section˜3.1. In Section 3.2, we provide a detailed description of the benchmark construction process.
@@ 3.1 Overview of MMSI-Bench
Table 1: Taxonomy of Multi-Image Spatial Reasoning.
Main Category Sub-Category Description Positional Relationship Camera-Camera Relationship between two camera viewpoints. Camera-Object Relationship between the camera and an object. Camera-Region Relationship between the camera and a semantic region (e.g., kitchen). Object-Object Relationship between two objects. Object-Region Relationship between an object and a semantic region. Region-Region Relationship between two different regions. Attribute Measurement Geometric properties (e.g., length, size) of elements. Appearance Visual characteristics (e.g., shape) of elements. Motion Camera The camera’s movement (e.g., direction). Object The movement of objects within the scene. Multi-Step Reasoning Complex tasks requiring a sequence of reasoning steps (e.g. navigation).
Humans possess the remarkable ability to infer spatial information about their environment by integrating visual inputs from different viewpoints or moments in time. We refer to this process as multi-image spatial reasoning, in which information from multiple images is combined to reconstruct the underlying scene and enable reasoning to answer complex spatial questions. To evaluate the multi-image spatial reasoning capability of multimodal large language models, we introduce MMSI-Bench. We further discuss additional motivations for multi-image input in Appendix C.
We categorize tasks based on three fundamental spatial elements: camera (the agent that captures the scenes), object (entities other than the agent within the scenes), and region (semantic areas within the scenes, such as a kitchen). Fundamental spatial understanding requires not only recognizing the positional relationships among these elements but also their attributes and potential motion. Guided by this taxonomy, we define ten spatial reasoning task types and one multi-step spatial reasoning category, which flexibly combines these basic types. The specific categories are presented in Table 1. Examples for all task categories are provided in Figure 2. Notably, our questions are designed to be human-answerable; thus, we do not include questions about camera attributes (parameters). Additionally, since regions are generally static, we do not include motion-related tasks for regions.
Figure 2: Representative MMSI-Bench samples from each category. Please zoom in to inspect image details. Questions and rationales are simplified for brevity; the complete versions appear in Appendix G. Correct answers are highlighted in green.
@@
In conclusion, MMSI-Bench consists of 1,000 multiple-choice question–answer pairs based on multiple images, each also accompanied by annotated reasoning processes explaining the answer. Every question is specifically designed such that it cannot be answered using a single image alone. Except for the Multi-Step Reasoning category, all other categories contain only two images per question. Detailed statistics on the proportion of questions in each category and other relevant information about MMSI-Bench are presented in Figure 3 and Table 3.1, respectively.
@@ 3.2 Benchmark Construction Process
Data collection. To comprehensively evaluate the capability of MLLMs across a variety of real-world scenarios, it is essential to curate data from diverse and representative sources. To ensure the complexity and authenticity of the evaluation, we exclusively utilize images captured from real-world scenes. Accordingly, we leverage images from a diverse collection of real-world datasets: indoor 3D scene datasets (Matterport3D (Chang et al., 2017) and ScanNet (Dai et al., 2017)), a local scene reconstruction dataset (DTU (Jensen et al., 2014)), autonomous driving datasets (nuScenes (Caesar et al., 2020), Waymo (Sun et al., 2020)), a robotics dataset (AgiBot-World (AgiBot-World-Contributors et al., 2025)), a video object segmentation dataset (DAVIS 2017 (Pont-Tuset et al., 2017)), and an egocentric video dataset (Ego4D (Grauman et al., 2022)). These diverse data sources enable the construction of various question-answer pairs, ensuring that our benchmark captures the complexity and diversity inherent in real-world environments.
@@
Question-answer design and reasoning process annotation. Using the collected datasets, six 3D vision researchers annotate QA pairs spanning the above eleven distinct task categories. Annotators are explicitly instructed to create a diverse set of questions that are as novel and as challenging as possible for MLLMs, aiming to push the boundaries of current model capabilities. At the same time, annotators are instructed to ensure that all questions are accurate, unambiguous, and answerable by humans. For each question, annotators begin by visually inspecting the image database to choose a set of relevant images. They then construct a question that cannot be answered based on any single image in isolation; rather, the question requires integrating information from multiple images in order to infer the actual scenario. To ensure diversity in question formats, no templates are used, and all questions are created in free form. Each question is designed as a four-option multiple-choice problem, with only one correct answer. The incorrect options are also carefully crafted to serve as plausible distractors. In addition, annotators provide detailed reasoning processes alongside each answer, which are used to facilitate filtering incorrect samples during quality control and to enable automated analysis of MLLMs’ reasoning processes, as described in Section 5.2.
@@
Quality control. To ensure the quality of the data, three additional reviewers, independent of the original annotators, systematically examine all data and remove any instances containing ambiguous questions (due to linguistic ambiguities or insufficient visual information), incorrect answers, or questions that can be answered using a single image or common sense alone. This rigorous quality control ensures our benchmark provides a reliable, high-quality evaluation. We annotate the difficulty level of the data based on human answering time; detailed information is provided in Appendix F.6.